OpenStudy (anonymous):

Factor the trinomial 2x^2 + 5x + 3. Write each factor as a polynomial in descending order

OpenStudy (anonymous):

u wnat afctorrr

OpenStudy (anonymous):

factors will be (2x+3) and (x+1)

OpenStudy (anonymous):

u deree

OpenStudy (jhannybean):

Follow this order. \[ 2x^2+5x +3\]\[\color{blue}2x^2 +5x+\color{blue}3\] You want to multiply these together to make your polynomial easier to factor.\[x^2 +5x+6\]3,2 add to give you 5, and multiply to give you 6.\[(x+3)(x+2)\]Since we multiplied by 2 in the beginning , we now have to divide by 2 to reinclude it in our equation. \[\left(x+\frac{3}{\color{blue}2}\right)\left(x+\frac{2}{\color{blue}{2}}\right)\]Finding out which fraction reduces, we finalize the last step by multiplying whatever DOESNT reduce diagonally over to the variable.\[(\color{blue}{2}x+3)(x+1)\]